# Towards Coordinate- and Dimension-Agnostic Machine Learning for Partial Differential Equations

### v1

@@ Appendix G Comparison between Different Coordinate Systems
Appendix H Initial Field Configurations for Testing
Here we list the initial field configurations used for testing in Section 4.1 (see Appendix H.1), Section 4.2 (see Appendix H.2), Section 4.2 (see Appendix H.3), and Section 4.3 (see Appendix H.4).
H.1 One-Dimensional Euclidean Space
In this case (Section 4.1), we work with the Patlak-Keller-Segel system. The initial field configurations The initial field configurations used in testing follow similar structure to Eq. (37), with:
@@ H.1 One-Dimensional Euclidean Space
We consider the same periodic space as in Appendix D for Patlak-Keller-Segel system, i.e. $x\in[-L,+L]$ with $L=10^{2}$ and a uniform grid of $\Delta x=1$. The transformation from this Cartesian coordinate $x$ to a non-Cartesian one $y$ follows Eq. (11). We use $y\in[-L,+L]$ with a uniform grid and a lattice spacing of $\Delta y=1$ when predicting the field evolution dynamics by integrating the trained neural network model from Section E.3 in this coordinate system.
We start with an initial field configuration that has not been used to generate the training data in Appendix D. The field $A$ is uniform everywhere and the field $B$ are combinations of localized configurations, as provided in Eq. (43). We create the field evolution dynamics using the timestep $\Delta t=10^{-1}$ for $t\in[0,5\times 10^{2}]$ by numerically integrating the PDEs Eq. (29).
H.2 Two-Dimensional Euclidean Space
In this case (Section 4.2), we work with the FitzHugh-Nagumo system. We start with the field configuration given by:
@@ H.2 Two-Dimensional Euclidean Space
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (6) for the total amount of time $\Delta T=8\times 10^{2}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section 4.2.
@@
H.3 Three-Dimensional Euclidean Space
In this case (Section 4.2), we work with the Barkley system. We start with the field configuration given by:
@@ H.3 Three-Dimensional Euclidean Space
Both $U$ and $V$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (28) for the total amount of time $\Delta T=2\times 10^{1}$ (so that the system has reached the stage of steady scroll wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section 4.2.
@@
H.4 A Two-Dimensional, Intrinsically Curved Surface
In this case (Section 4.3), we work with the FitzHugh-Nagumo system. On the spherical surface, the FitzHugh-Nagumo system (Section 3.1) written in the $(\theta,\phi)$-coordinate can be described by the following PDEs:
$\begin{split}\partial_{t}V(\theta,\phi,t)&=\frac{1}{R^{2}}\left(\cot\theta\partial_{\theta}+\partial_{\theta}^{2}+\frac{1}{\sin^{2}\theta}\partial_{\phi}^{2}\right)V(\theta,\phi,t)\\ &\ \ \ \ \ \ \ \ +V(\theta,\phi,t)-\frac{1}{3}V^{3}(\theta,\phi,t)-W(x,t)\ ,\\ \partial_{t}W(\theta,\phi,t)&=\varepsilon\left[V(\theta,\phi,t)+\beta-\gamma W(\theta,\phi,t)\right]\ .\end{split}$ (48)
@@ H.4 A Two-Dimensional, Intrinsically Curved Surface
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (48) for the total amount of time $\Delta T=2\times 10^{2}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section 4.3.
@@
$\left(\Delta\rho,\Delta\varphi\right)=\left(\frac{\rho_{\max}-\rho_{\max}}{300},\frac{\varphi_{\max}-\varphi_{\max}}{200}\right)\ .$ (51)
@@
$\left(\Delta\theta,\Delta\phi\right)=\left(\frac{\theta_{\max}-\theta_{\max}}{300},\frac{\phi_{\max}-\phi_{\max}}{200}\right)\ .$ (52)
@@
Appendix I Geographic and Stereographic Coordinates
@@
We aim to demonstrate that, using the neural network model trained in one dimension in Sec. D, we can predict the correct field evolution dynamics on the surface of the sphere regardless of the chosen coordinate system. In other words, not only do we need to generate a prediction, but also compare it with the expected true dynamics. We begin with an initial condition defined in $(\theta,\phi)$-space with uniform $(\Delta\theta,\Delta\phi)$ discretization. We then compute the field evolution in the $(\theta,\phi)$-coordinate using direct integration of the true PDE. Next, we determine by extrapolation the corresponding initial condition in $(\rho,\varphi)$-space with uniform $(\Delta\rho,\Delta\varphi)$ discretization and use it to generate the field evolution in $(\rho,\varphi)$-coordinates via the trained neural network model in Section D. To make comparison, we transform both the generated datasets to the three-dimensional Euclidean $(x,y,z)$-space. These steps are similar to what has been demonstrated for one-dimensional case, see Fig. 6. For our spatially liberated neural network model, all these generated datasets are expected to be in agreement, which we confirm through our results.
Appendix J Comparing “True” and Learned Dynamics
@@
J.1 1-, 2- and 3-Dimensional Euclidean Space
@@
[…]
@@
J.2 2-Dimensional Intrinsically curved Surface
@@
[…]
Figure 8: Comparisons between “true” and neural learned dynamics in two-dimensional intrinsically curved surface embedded in three-dimensional Euclidean space. Here we study FitzHugh-Nagumo system, where the deviation range for each field is equal to its dynamical range, with negative deviation colored red-ish and positive deviation colored blue-ish. We show the deviation between the top and bottom rows of Fig. 4D1 in (A), of Fig. 4D2 in (B), of Fig. 4D3 in (C), and of Fig. 4D4 in (D). Note that, for brevity, we only label the first plot, since the rest are identical.

### v2

@@ Appendix H Comparison between Different Coordinate Systems
Appendix I Initial Field Configurations for Testing
Here we list the initial field configurations used for testing in Section D.1 (see Appendix I.1), Section D.2 (see Appendix I.2), Section D.2 (see Appendix I.3), and Section D.3 (see Appendix I.4).
I.1 One-Dimensional Euclidean Space
In this case (Section D.1), we work with the Patlak-Keller-Segel system. The initial field configurations The initial field configurations used in testing follow similar structure to Eq. (38), with:
@@ I.1 One-Dimensional Euclidean Space
We consider the same periodic space as in Appendix E for Patlak-Keller-Segel system, i.e. $x\in[-L,+L]$ with $L=10^{2}$ and a uniform grid of $\Delta x=1$. The transformation from this Cartesian coordinate $x$ to a non-Cartesian one $y$ follows Eq. (25). We use $y\in[-L,+L]$ with a uniform grid and a lattice spacing of $\Delta y=1$ when predicting the field evolution dynamics by integrating the trained neural network model from Section F.3 in this coordinate system.
We start with an initial field configuration that has not been used to generate the training data in Appendix E. The field $A$ is uniform everywhere and the field $B$ are combinations of localized configurations, as provided in Eq. (44). We create the field evolution dynamics using the timestep $\Delta t=10^{-1}$ for $t\in[0,5\times 10^{2}]$ by numerically integrating the PDEs Eq. (6).
I.2 Two-Dimensional Euclidean Space
In this case (Section D.2), we work with the FitzHugh-Nagumo system. We start with the field configuration given by:
@@ I.2 Two-Dimensional Euclidean Space
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (4) for the total amount of time $\Delta T=8\times 10^{2}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.2.
@@
I.3 Three-Dimensional Euclidean Space
In this case (Section D.2), we work with the Barkley system. We start with the field configuration given by:
@@ I.3 Three-Dimensional Euclidean Space
Both $U$ and $V$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (5) for the total amount of time $\Delta T=2\times 10^{1}$ (so that the system has reached the stage of steady scroll wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.2.
@@
I.4 A Two-Dimensional, Intrinsically Curved Surface
In this case (Section D.3), we work with the FitzHugh-Nagumo system. On the spherical surface, the FitzHugh-Nagumo system written in the $(\theta,\phi)$-coordinate can be described by the following PDEs:
$\begin{split}\partial_{t}V(\theta,\phi,t)&=\frac{1}{R^{2}}\left(\cot\theta\partial_{\theta}+\partial_{\theta}^{2}+\frac{1}{\sin^{2}\theta}\partial_{\phi}^{2}\right)V(\theta,\phi,t)\\ &\ \ \ \ \ \ \ \ +V(\theta,\phi,t)-\frac{1}{3}V^{3}(\theta,\phi,t)-W(\theta,\phi,t)\ ,\\ \partial_{t}W(\theta,\phi,t)&=\varepsilon\left[V(\theta,\phi,t)+\beta-\gamma W(\theta,\phi,t)\right]\ .\end{split}$ (49)
@@ I.4 A Two-Dimensional, Intrinsically Curved Surface
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (49) for the total amount of time $\Delta T=2\times 10^{2}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.3.
@@
$\left(\Delta\rho,\Delta\varphi\right)=\left(\frac{\rho_{\max}-\rho_{\min}}{300},\frac{\varphi_{\max}-\varphi_{\min}}{200}\right)\ .$ (52)
@@
$\left(\Delta\theta,\Delta\phi\right)=\left(\frac{\theta_{\max}-\theta_{\min}}{300},\frac{\phi_{\max}-\phi_{\min}}{200}\right)\ .$ (53)
@@
I.5 A Two-Dimensional, Flat Surface, with Non-Trivial Topology
We work with the FitzHugh-Nagumo system for this demonstration (Section D.4). In 2-D polar coordinate system, the FitzHugh-Nagumo PDEs can be expressed in the $(\rho,\theta)$-coordinate as:
$\begin{split}\partial_{t}V(\rho,\theta,t)&=\frac{1}{\rho}\partial_{\rho}\left[\rho\partial_{\rho}V(\rho,\theta,t)\right]+\frac{1}{\rho^{2}}\partial_{\theta}^{2}V(\rho,\theta,t)\\ &\ \ \ \ \ \ \ \ +V(\rho,\theta,t)-\frac{1}{3}V^{3}(\rho,\theta,t)-W(\rho,\theta,t)\ ,\\ \partial_{t}W(\rho,\theta,t)&=\varepsilon\left[V(\theta,\phi,t)+\beta-\gamma W(\theta,\phi,t)\right]\ .\end{split}$ (54)
We start with the field configuration given by:
$\begin{split}V(\rho,\theta,0)&=V_{01}\exp\left[-\frac{(\rho-\rho_{V1})^{2}}{2\sigma_{\rho V1}}-\frac{(\theta-\theta_{V1})^{2}}{2\sigma_{\theta V1}}\right]\\ &+V_{02}\exp\left[-\frac{(\rho-\rho_{V2})^{2}}{2\sigma_{\rho V2}}-\frac{(\theta-\theta_{V2})^{2}}{2\sigma_{\theta V2}}\right]\\ &+V_{01}\exp\left[-\frac{(\rho-\rho_{V1})^{2}}{2\sigma_{\rho V1}}-\frac{(\theta-\theta_{V1}-\pi/2)^{2}}{2\sigma_{\theta V1}}\right]\\ &+V_{02}\exp\left[-\frac{(\rho-\rho_{V2})^{2}}{2\sigma_{\rho V2}}-\frac{(\theta-\theta_{V2}-3\pi/2)^{2}}{2\sigma_{\theta V2}}\right]\ ,\\ W(\rho,\varphi,0)&=W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\theta-\theta_{W})^{2}}{2\sigma_{\theta W}}\right]\\ &+W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\theta-\theta_{W}-\pi/2)^{2}}{2\sigma_{\theta W}}\right]\\ &+W_{0}\exp\left[-\frac{(\rho-\rho_{W})^{2}}{2\sigma_{\rho W}}-\frac{(\theta-\theta_{W}-3\pi/2)^{2}}{2\sigma_{\theta W}}\right]\ .\end{split}$ (55)
where:
$\begin{split}(V_{01},\rho_{V1},\theta_{V1},\sigma_{\rho V1},\sigma_{\theta V1})&=(1,140,\pi/6,15,0.15)\ ,\\ (V_{02},\rho_{V2},\theta_{V2},\sigma_{\rho V2},\sigma_{\theta V2})&=(1,160,\pi/4,10,0.1)\ ,\\ (W_{0},\rho_{W},\theta_{W},\sigma_{\rho W},\sigma_{\theta W})&=(1,150,\pi/4,10,0.2)\ .\end{split}$ (56)
Both $V$ and $W$ are combinations of localized configurations. Then, we numerically integrate the PDEs Eq. (54) for the total amount of time $\Delta T=10^{3}$ (so that the system has reached the stage of steady spiral wave generation) from the above configurations. We use the final-time configuration as the initial field configuration for the study in Section D.4.
We work on the annulus confined between the radius value $\rho\in[\rho_{\min},\rho_{\max}]$, where $\rho_{\min}=10^{2}$, $\rho_{\max}=2\times 10^{2}$. We discretize the coordinates $(\rho,\theta)$ uniformly with the lattice-spacing
$\left(\Delta\rho,\Delta\theta\right)=\left(1,2\pi/800\right)\ .$ (57)
Then, we compute the field evolution dynamics on the spherical surface with the true dynamics and with the neural network learned dynamics, using the timestep $\Delta t=10^{-1}$ for $t\in[0,5\times 10^{1}]$.
Appendix J Geographic and Stereographic Coordinates
@@ I.5 A Two-Dimensional, Flat Surface, with Non-Trivial Topology
We aim to demonstrate that, using the neural network model trained in one dimension in Sec. E, we can predict the correct field evolution dynamics on the surface of the sphere regardless of the chosen coordinate system. In other words, not only do we need to generate a prediction, but also compare it with the expected true dynamics. We begin with an initial condition defined in $(\theta,\phi)$-space with uniform $(\Delta\theta,\Delta\phi)$ discretization. We then compute the field evolution in the $(\theta,\phi)$-coordinate using direct integration of the true PDE. Next, we determine by extrapolation the corresponding initial condition in $(\rho,\varphi)$-space with uniform $(\Delta\rho,\Delta\varphi)$ discretization and use it to generate the field evolution in $(\rho,\varphi)$-coordinates via the trained neural network model in Section E. To make comparison, we transform both the generated datasets to the three-dimensional Euclidean $(x,y,z)$-space. These steps are similar to what has been demonstrated for one-dimensional case, see Fig. 9. For our spatially liberated neural network model, all these generated datasets are expected to be in agreement, which we confirm through our results.
Appendix K Comparing “True” and Learned Dynamics
@@
K.1 1-, 2- and 3-Dimensional Euclidean Space
@@
  • Fig. 10A1: $_{A}=1.67\times 10^{-5}, MSE$$_{B}=4.90\times 10^{-6}$.
  • Fig. 10A2: $_{A}=8.56\times 10^{-6}, MSE$$_{B}=7.76\times 10^{-5}$.
  • Fig. 10A3: $_{A}=1.24\times 10^{-5}, MSE$$_{B}=7.98\times 10^{-5}$.
  • Fig. 10B1: $_{V}=2.01\times 10^{-3}, MSE$$_{W}=1.20\times 10^{-5}$.
  • Fig. 10B2: $_{V}=1.62\times 10^{-2}, MSE$$_{W}=1.49\times 10^{-4}$.
  • Fig. 10B3: $_{V}=3.03\times 10^{-2}, MSE$$_{W}=9.16\times 10^{-4}$.
  • Fig. 10B4: $_{V}=4.67\times 10^{-2}, MSE$$_{W}=1.39\times 10^{-3}$.
  • Fig. 10C1: $_{U}=0, MSE$$_{V}=0$ (initial field configurations are identical).
  • Fig. 10C2: $_{U}=3.19\times 10^{-3}, MSE$$_{V}=3.35\times 10^{-4}$.
  • Fig. 10C3: $_{U}=6.24\times 10^{-3}, MSE$$_{V}=7.44\times 10^{-4}$.
  • Fig. 10C4: $_{U}=9.31\times 10^{-3}, MSE$$_{V}=1.15\times 10^{-3}$.
@@
K.2 2-Dimensional Intrinsically curved Surface
@@
  • Fig. 11A: $_{V}=5.85\times 10^{-5}, MSE$$_{W}=1.35\times 10^{-8}$.
  • Fig. 11B: $_{V}=1.36\times 10^{-3}, MSE$$_{W}=1.51\times 10^{-5}$.
  • Fig. 11C: $_{V}=2.34\times 10^{-3}, MSE$$_{W}=6.56\times 10^{-5}$.
  • Fig. 11D: $_{V}=4.47\times 10^{-3}, MSE$$_{W}=1.01\times 10^{-4}$.
Figure 11: Comparisons between “true” and neural learned dynamics in two-dimensional intrinsically curved surface embedded in three-dimensional Euclidean space. Here we study FitzHugh-Nagumo system, where the deviation range for each field is equal to $10\%$ of its dynamical range, with negative deviation colored red-ish and positive deviation colored blue-ish. We show the deviation between the top and bottom rows of Fig. 6D1 in (A), of Fig. 6D2 in (B), of Fig. 6D3 in (C), and of Fig. 6D4 in (D). Note that, for brevity, we only label the first plot, since the rest are identical.
K.3 2-Dimensional Non-Trivial Topology
We show the deviation for FitzHugh-Nagumo system in two-dimensional flat space with an annulus topology, in Fig. 12.
@@
  • Fig. 12A: $_{V}=0, MSE$$_{W}=0$ (initial field configurations are identical).
  • Fig. 12B: $_{V}=9.25\times 10^{-4}, MSE$$_{W}=1.11\times 10^{-5}$.
  • Fig. 12C: $_{V}=1.64\times 10^{-3}, MSE$$_{W}=4.69\times 10^{-5}$.
  • Fig. 12D: $_{V}=3.33\times 10^{-3}, MSE$$_{W}=7.54\times 10^{-5}$.
Figure 12: Comparisons between “true” and neural learned dynamics in two-dimensional flat surface embedded with an annulus topology. Here we study FitzHugh-Nagumo system, where the deviation range for each field is equal to $10\%$ of its dynamical range, with negative deviation colored red-ish and positive deviation colored blue-ish. We show the deviation between the top and bottom rows of Fig. 7A in (A), of Fig. 7B in (B), of Fig. 7C in (C), and of Fig. 7D in (D).
Appendix L On the Quality of the Prediction using Noisy Training Data
In fact, our study was developed with real experiments in mind: we are preparing to apply it to a real bacteria-chemical system as reported in Phan et al. (2024), which is also the work that proposed the modified Patlak-Keller-Segel PDE used in the main text. The model in Eq. (24) is presented in dimensionless form; Table 1 summarizes the mapping between dimensional parameters and their dimensionless counterparts.
Asp dynamics Bacteria dynamics Non-Dimensionalized
@@
Table 1: The best-estimated parameters are those reported in Phan et al. (2024), based on an analysis of traveling-wave experiments with E. coli strain HE205 Cremer et al. (2019). Unlisted non-dimensionalized parameters are, by convention, implied to be set to $1$, i.e. $D_{A}=1$, $\gamma_{0}=1$, $B_{C}=1$, and $A_{h}=1$. In those experiments, the maximum Asp concentration $\approx 100\mu\mathrm{M}$ and the maximum cell density $\approx 6\mathrm{OD}$, which correspond to the nondimensional maxima $A_{\max}\approx 27$ and $B_{\max}\approx 1$. The non-dimensionalized unit-time and unit-length correspond to the time-scale $\tau\approx 1.4\mathrm{s}$ and the length-scale $\lambda\approx 33\mu\mathrm{m}$. While these values can vary across strains, they are expected to remain within the same order of magnitude.
We can show that the experimental noise for this system is relatively low, can be made much lower, and also can be estimated as follows. The microfluidic channel has width $W\approx 1.2\mathrm{mm}$ and depth $H\approx 100\mu\mathrm{m}$ (which can be increased). The bacterial density in the traveling wave is around $n\sim 2\mathrm{OD}$ (with $1\mathrm{OD}\approx 8\times 10^{8}\mathrm{cells/mL}$), and wave peaks can reach $\sim 5\mathrm{OD}$. The effective one-dimensional bin size in our discretization is $\Delta x=1$ (dimensionless), corresponding to $\approx 33\mu\mathrm{m}$ physically (see the caption of Table 1). Hence, the expected number of bacteria per bin is
$N\sim nWH\Delta x\approx 8\times 10^{3}\mathrm{cells}\ .$
Assuming purely Poisson (counting) statistics, which approach a Gaussian in the large-number limit, the relative shot noise in the bacterial density field $B(x,t)$ is $\sigma\sim N^{-1/2}\sim 1\%$. Using a deeper device, e.g. increasing $H$ to $1.2\mathrm{mm}$, would reduce this further to
$\sigma\approx 0.3\%\ .$
The chemical concentration field $A(x,t)$ is even less noisy (effectively negligible in experiments), owing to its much higher molecular counts and diffusivity compared to those of bacteria. We believe these empirical conditions make the noise-free assumption acceptable here, while still laying groundwork for further studies.
If we add this $\sigma\sim 0.3\%$ noise to the training data and apply Savitzky-Golay smoothing before computing derivatives, the neural network – after the same training mentioned in Appendix E.3 – can still make reasonable and correct-qualitative predictions (especially in the region where both bacteria and chemical field values are high, where noise does not dominate training, which is from $t=0$ to $250$) for the initial condition as used in Fig. 2B1 and Fig. 2B2 of the main manuscript. Here we report the MSE for the values between $t=0$ and $t=250$ (see Fig. 13A):
  • $_{A}=4.76\times 10^{-3}, MSE$$_{B}=1.24\times 10^{-3}$.
Despite a substantially higher MSE after adding noise, the practical implication does not change: the prediction correctly tracks the traveling-wave front (speed, shape, persistence). Also, it should be noted that our current noise‐handling strategy is simple – simply applying a mild smoothing filter before training – but already yields acceptable predictions.
Figure 13: The predictions for Patlak-Keller-Segel field evolution dynamics in a periodic one-dimensional space, using training data corrupted by realistic noise. The initial field configurations are the same as used in Fig. 2B1 and Fig. 2B2. The gray boxes capture the spacetime regions between $t=0$ and $t=200$. (A1) The time progression of the fields when using the neural network trained the same way as in Appendix E.3. (A2) The field values difference between the neural network predicted Fig. 13A1 and the true Fig. 2B1. (B1) The time progression of the fields when using the neural network trained with $L^{1}$-norm (instead of $L^{2}$-norm). (B2) The field values difference between the neural network predicted Fig. 13B1 and the true Fig. 2B1.
There are several ways to improve performance; for instance, changing the training procedure by switching from an $L^{2}$-norm to a $L^{1}$-norm for the loss function makes the neural network model much more robust to noise. For comparison, we report this MSE between $t=0$ and $t=250$:
  • $_{A}=5.26\times 10^{-4}, MSE$$_{B}=1.32\times 10^{-4}$.
These results are now smaller by more than an order of magnitude and could be reduced further with more sophisticated noise‐resistant strategies.
Appendix M General PDE Example
The PCA model was fit to the 200 numerical trajectories of the 1-D harmonic oscillator. For these solutions, we approximate the first and second derivatives of $u$ using finite differences on a set of 200 solutions (with random initial conditions), and perform PCA, we obtained explained variance ratios of 6.7e1,3.3e1,6.4e66.7e-1,3.3e-1,6.4e-6, denoting the 2-dimensional linear structure of the manifold.
After identifying the operator, we proceed to train a PINN in the manner discussed in the main text. The PINN architecture consists of three layers with width 64 and a $\tanh$ activation function, trained using Adam Kingma (2014). The mean square error between a classical PINN solution and the coordinate-free version visualized in Fig. 14 is approximately $4e^{-4}$, while the square difference between the these solutions is depicted the right sub-plot of Figure 14.
Figure 14: Left: Solution of the Helmholtz equation obtained using classical PINN training. Middle: Solution of the Helmholtz equation obtained using a PINN with a coordinate-free PDE loss, where the PDE operator is identified from trajectories of the 1-D harmonic oscillator (see text for details). Right: Squared error between classical PINN solution and coordinate-free PINN solution to the 2D helmholtz problem.
Appendix N Vector PDE Example
The solution to the flow around a cylinder example is computed on the unit square, $(x,y)\in[0,1]^{2}\subset\mathbb{R}^{2}$, discretized using a regular grid of $100\times 100$ points. A cylinder of radius $r=0.1$ is centered at the center of the domain. The velocity field $v(x,y,t)$ is set initially to zero $v(x,y,0)=(0,0)$, with the exception of the left boundary of the square where we set $v(0,x,t)=1$. The viscosity $\nu=\eta/\rho$ is set to $\nu=0.01$.
After obtaining a numerical solution $\hat{v}$ using the setup described above, we compute the corresponding vector features $\quantity{(v\cdot\gradient)v,\Delta v,\gradient p}$. Using least squares, we determine the coefficients that best approximate $\partialderivative{v}{t}$ as a linear combination of these features, which for the example presented in Figure 5 are $(1.01,0.01,1.00)$ (rounded to decimal digits).
Using these coefficients, we integrate the momentum equation for the Taylor-Green vortex problem described below, using an Euler integration scheme, while assuming that the true pressure field is known for all time. We integrate on the domain $0,2\pi]^{2}\subset\mathbb{R}^{2}$, discretized using a $100\times 100$ grid. The solution depicted in the bottom of Figure 5 is obtained after 10 Euler time steps of size $dt=0.01$, with a relative $\ell^{2}$ error of 1.2e-1 (with respect to the analytical solution).
The Taylor-Green Vortex dynamics (Taylor & Green, 1937) are a special case of the incompressible Navier-Stokes equation, defined as follows:
For the (periodic) domain ${(x,y)}\in[0,2\pi]^{2}$, we consider the initial velocity field to take the form
$v_{1}=\cos(x)\sin(y),\quad v_{2}=-\sin(x)\cos(y)$ (64)
In this case, it is known that the velocity field as a function of time takes the form:
$v(x,y,t)$ $=v(x,y,0)e^{-2\frac{\eta}{\rho}t}$ (65)
$p(x,y,t)$ $=\frac{\rho}{4}\quantity({\cos(2x)+\cos(2y)})e^{-4\frac{\eta}{\rho}t}$ (66)
To further simplify the PDE identification and integration problem, we assume that the scalar field $p(x,y,t)$ is known at all times, and that the continuity equation is also known and does not need to be identified. In that way, we have reduced the problem to identifying a time-evolution vector-PDE (i.e. the momentum equation) for which we can use a simple Euler scheme to integrate (in order to verify our coordinate-free learning procedure).